OpenStudy (anonymous):

Will give medal/fan/etc. Three cards are dealt from a well-shuffled standard deck. f. Find the probability that either one or two of the

OpenStudy (anonymous):

I've worked out part of it, and i know the answer, but I don't know why my answer doesn't match the given answer.

OpenStudy (anonymous):

1 Ace = 4/52 * 48/51 * 47/50 * 3 3 because if A = Ace and X= Not ace, the combos for three cards are AXX, XAX, XXA

OpenStudy (anonymous):

2 Ace = 4/52 * 3/51 * 48/50 * 3 3 because if A = Ace and X = not ace, the combos for the there cards are AAX, AXA, XAA

OpenStudy (anonymous):

Given answer = 48/221

OpenStudy (anonymous):

I thought that I had to add them together because "either" means that its mutually exclusive

OpenStudy (anonymous):

That will mean that whatever I add to has to add up to 173/221 or 4325/5525

OpenStudy (mathmath333):

first see if ur question is complete

OpenStudy (anonymous):

Well the question has parts a-f, but I finished a-e